# LancBiO: Dynamic Lanczos-aided Bilevel Optimization via Krylov Subspace

### v1

@@ Appendix E Proof of Smoothness of $y^{*}$ and $\varphi$
By observing that $\left(\mu_{s_{1}}^{\left(n_{1}\right)},y_{s_{1}}^{\left(n_{1}\right)}\right)$ and $\left(y_{s_{1}}^{\left(n_{1}\right)}\right)^{\top}R_{n_{1}}y_{s_{1}}^{\left(n_{1}\right)}$ can also be categorized into Case I or Case II, repeat this process and construct a sequence $\{(s_{t},n_{t})\}_{t=0}^{l+1}$ with $1\leq n_{l+1}<n_{l}<\cdots<n_{1}<n_{0}=j$ until
$\left(y_{s_{l+1}}^{\left(n_{l+1}\right)}\right)^{\top}R_{j}y_{s_{l+1}}^{\left(n_{l+1}\right)}<\frac{3}{8}-\frac{\varepsilon_{n_{l+1}}}{2}.$
Inequality (42) holds when the superscript $j=1$ since $T_{1}=[\alpha_{1}],\ y_{1}^{(1)}=1$ and $z^{(1)}_{1}=q_{1}$. Therefore, we will build up $\{(s_{t},n_{t})\}_{t=0}^{l+1}$ in finite steps, resulting in the following estimation.
@@
Any Rayleigh quotient of $A_{j}^{*}$ is bounded by its eigenvalues (Parlett, 1998), i.e., for any $\nu^{(j)}$,
@@
where $\lambda_{min}^{(j)}$ and $\lambda_{max}^{(j)}$ are the minimum and maximal eigenvalue of $A_{j}^{*}$, respectively. Based on 3.2 and 3.3, we thus complete the proof. ∎
@@ G.1 Proof sketch
The proof of Lemma 3.12 is structured by four principal steps.
Step1: extending $\varepsilon_{j}$ to $\tilde{\varepsilon}_{j}$ within the lemmas detailed in Appendix F.
@@
Note that we can also view the $A_{1},b_{1}$ as reference values. In this way, we denote the similar quantities
@@ G.2 Extended lemmas from Appendix F
(Extended version of Lemma F.2) Suppose Assumptions 3.1 to 3.3 hold. The dynamic Lanczos process in Algorithm 2 with normalized $q_{1}$ and $\alpha_{j},\beta_{j},q_{j}$ satisfies
@@ G.3 Proof of Step2
where $T_{j}$ is nonsingular because of Lemma G.2. By (43), (46), and (49),
@@
In this way, $\left|\beta_{j+1}e_{j}^{\top}\left(T_{j}\right)^{-1}e_{1}\right|$ can be seen as the residual in the $j$-th step of the classic Lanczos process with the positive-definite matrix $T$ and the iniitial vector $e_{1}$. Since the eigenvalues of $T$ satisfy (52), it follows from the standard convergence property of the Lanczos process (Greenbaum, 1997) that
@@
G.4 Proof details of $\mu_{g}-\frac{2\sqrt{3}}{3}\left(j+1\right)^{3}L_{gyy}{\varepsilon}_{j}>0$
In this part, we will give the detailed proof of $\mu_{g}-\frac{2\sqrt{3}}{3}\left(j+1\right)^{3}L_{gyy}{\varepsilon}_{j}>0$. At the same time, we demonstrate that, with appropriate step sizes, the auxiliary variable $v_{k}$ is bounded, thereby ensuring that the hyper-gradient estimator (5) remains bounded, which ensures the stable behavior of the dynamic Lanczos process,
@@ G.4 Proof details of $\mu_{g}-\frac{2\sqrt{3}}{3}\left(j+1\right)^{3}L_{gyy}{\varepsilon}_{j}>0$
Suppose Assumptions 3.1, 3.2, 3.3, 3.6, 3.9, 3.10 and 3.11 hold. If within each epoch, we set the step size $\theta\sim\mathcal{O}(\frac{1}{m^{4}})$ a constant for $y$ and the step size for $x$ as zero in the first $m_{0}\sim\mathcal{O}(1)$ steps, and the others as an appropriate constant $\lambda\sim\mathcal{O}(\frac{1}{m^{4}})$, then for any epoch,
@@
Then, we will prove by induction. At the beginning of the algorithm, the condition (53) reveals
@@
Then by setting the stepsize
$\lambda\leq\frac{\sqrt{3}\tilde{\varepsilon}}{4(m+1)^{4}L_{gyy}}\left(C_{fx}+\frac{1}{\mu_{g}^{2}}\left(3\sqrt{\tilde{\kappa}}C_{r}+C_{fx}\right)\right)^{-1},$
$\theta\leq\frac{\sqrt{3}\tilde{\varepsilon}}{4(m+1)^{4}L_{gyy}}C_{gy}^{-1},$
@@
$\tilde{\varepsilon}_{j+1}\leq\lambda\sum_{i=1}^{i=j}{\left\|\widetilde{\nabla}\varphi_{i}\right\|}+\theta\sum_{i=1}^{i=j}{\left\|\nabla_{y}g\left(x_{i+1},y_{i}\right)\right\|}\leq\frac{\sqrt{3}\tilde{\varepsilon}}{2(m+1)^{3}L_{gyy}}.$ (54)
@@
Set
$\lambda\leq\frac{\sqrt{3}\mu_{g}}{8(m+1)^{4}L_{gyy}}\left(C_{fx}+\frac{1}{\mu_{g}^{2}}\left(3\sqrt{\tilde{\kappa}}C_{r}+C_{fx}\right)\right)^{-1}\left(1+\frac{L_{gx}}{\mu_{g}}+\left(1-\theta\mu_{g}\right)^{\frac{1}{2}}\left(\frac{L_{gx}}{\mu_{g}}\right)\right)^{-1},$
it yields from (11), (56), (59) that
@@
As for the next epoch, denoting $C_{v}=\frac{1}{\mu_{g}}\left(3\sqrt{\tilde{\kappa}}C_{r}+C_{fx}\right)$ we have
@@
by setting $m,\tilde{\varepsilon}$ such that
@@
Suppose Assumptions 3.1, 3.2, 3.3, 3.6, 3.9, 3.10 and 3.11 hold. If within each epoch, we set the step size $\theta\sim\mathcal{O}(\frac{1}{m^{4}})$ a constant for $y$ and the step size for $x$ as zero in the first $m_{0}\sim\mathcal{O}(1)$ steps, and the others as an appropriate constant $\lambda\sim\mathcal{O}(\frac{1}{m^{4}})$, then we have the following.
@@
Appendix H Proof of The Main Theorems
In this section, we provide proof of the main theorem presented in Section 3.2. Let $A_{k}=\nabla_{yy}g(x_{k},y_{k})$ and $b_{k}=\nabla_{y}f(x_{k},y_{k})$, and let the reference values be $A_{k}^{*}=\nabla_{yy}^{2}g\left(x_{k},y^{*}_{k}\right)$, $b^{*}_{k}=\nabla_{y}f\left(x_{k},y^{*}_{k}\right)$ and $v_{k}^{*}=(A_{k}^{*})^{-1}b_{k}^{*}$.
The following lemma displays the descent properties of the iterates $\{y_{k}\}$.
@@ Appendix H Proof of The Main Theorems
Suppose Assumptions 3.2 and 3.3 hold, we have
@@
Algorithm 2 executes a single-step gradient descent on the strongly convex function $g(x_{k+1},\cdot)$ during the outer iteration. Leveraging the established convergence properties of strongly convex functions (Nesterov et al., 2018), we are thus able to derive the following.
@@
Suppose Assumptions 3.1, 3.2, 3.3, 3.6, 3.10 and 3.11 hold. Within each epoch, we set the step size $\theta\sim\mathcal{O}(\frac{1}{m^{4}})$ a constant for $y$ and the step size for $x$ as zero in the first $m_{0}$ steps, and the others as an appropriate constant $\lambda\sim\mathcal{O}(\frac{1}{m^{4}})$, then the iterates
@@
~$\left\|\widetilde{\nabla}\varphi\left(x_{k}\right)-\nabla\varphi\left(x_{k}\right)\right\|^{2}\leq$ $~{}3L_{gx}^{2}\iota^{2\left(j-m_{0}\right)+h\left(m-m_{0}\right)}\delta_{0}+18\left(1-\gamma\right)^{-1}\lambda^{2}L_{gx}^{2}m^{2}\omega_{\varphi}\sum_{t=0}^{j-1}{\left\|\nabla\varphi\left(x_{mh+t}\right)\right\|^{2}}$
$+18\left(1-\gamma\right)^{-1}\lambda^{2}L_{gx}^{2}m^{2}\omega_{\varphi}\sum_{e=0}^{h-1}{\iota^{2\left(h-e\right)\left(m-m_{0}\right)}\sum_{t=0}^{m-1}{\left\|\nabla\varphi\left(x_{me+t}\right)\right\|^{2}}},$
where $10<\gamma<$ 1, $m_{0}\sim\mathcal{O}(\log m)$ are constants, $m$ is the subspace dimension, and $\delta_{0}$ is a constant defined in the proof.
@@
By (H), conclusion from Lemma G.3 and the Young’s inequality,
@@
Now, we define
$\delta_{k}=\left(\frac{L_{fx}^{2}\mu_{g}^{2}+L_{gxy}^{2}C_{fg}^{2}}{L_{gx}^{2}\mu_{g}^{2}}\right)\left\|y_{k}-y_{k}^{*}\right\|^{2}+\left\|v_{k}-v_{k}^{*}\right\|^{2},$
and add $\left(\frac{L_{fx}^{2}\mu_{g}^{2}+L_{gxy}^{2}C_{fg}^{2}}{L_{gx}^{2}\mu_{g}^{2}}\right)\left\|y_{k}-y_{k}^{*}\right\|^{2}$ on both sides of (H). Combine it with Lemma H.1,
$\delta_{k}\leq$ $~{}\omega_{\varepsilon}\left(\varepsilon_{j}^{h}\right)^{2}+\frac{L_{gy}^{2}}{\mu_{g}^{2}}\tilde{\kappa}\left(j\right)\left(\frac{\sqrt{\tilde{\kappa}\left(j\right)}-1}{\sqrt{\tilde{\kappa}\left(j\right)}+1}\right)^{2j}\left\|v_{mh}-v_{mh}^{*}\right\|^{2}$
$+\frac{\tilde{\kappa}\left(j\right)}{\mu_{g}^{2}}\left(\frac{\mu_{g}L_{fy}+C_{fy}L_{gyy}}{\mu_{g}}\left(1+\frac{L_{gx}}{\mu_{g}}\right)\right)^{2}\left(\frac{\sqrt{\tilde{\kappa}\left(j\right)}-1}{\sqrt{\tilde{\kappa}\left(j\right)}+1}\right)^{2j}\left\|x_{mh+1}-x_{mh}\right\|^{2}$
$+\left(1+\sigma\right)^{j}\left(1-\theta\mu_{g}\right)^{j}\frac{L_{fx}^{2}\mu_{g}^{2}+L_{gxy}^{2}C_{fg}^{2}}{L_{gx}^{2}\mu_{g}^{2}}\left\|y_{mh}-y_{mh}^{*}\right\|^{2}$
$+\left(1+\frac{1}{\sigma}\right)\left(1-\theta\mu_{g}\right)\frac{L_{fx}^{2}\mu_{g}^{2}+L_{gxy}^{2}C_{fg}^{2}}{L_{gx}^{2}\mu_{g}^{2}}\left(\frac{L_{gx}}{\mu_{g}}\right)^{2}\sum_{r=0}^{j}{\left(1+\sigma\right)^{r}\left(1-\theta\mu_{g}\right)^{r}\left\|x_{k-r}-x_{k-r-1}\right\|^{2}}.$ (66)
We set $\sigma,\theta$ to satisfy $\left(1+\sigma\right)\left(1-\theta\mu_{g}\right)<1$, and since Lemma G.4 reveals that under the appropriate step-size setting $\lambda\sim\mathcal{O}(\frac{1}{m^{4}})$ and $\theta\sim\mathcal{O}(\frac{1}{m^{4}})$, $\tilde{\kappa}(j)\leq\tilde{\kappa}$ for some $\tilde{\kappa}>0$, we define

### v2

@@ Appendix E Proof of Smoothness of $y^{*}$ and $\varphi$
By observing that $\left(\mu_{s_{1}}^{\left(n_{1}\right)},y_{s_{1}}^{\left(n_{1}\right)}\right)$ and $\left(y_{s_{1}}^{\left(n_{1}\right)}\right)^{\top}R_{n_{1}}y_{s_{1}}^{\left(n_{1}\right)}$ can also be categorized into Case I or Case II, we can repeat this process and construct a sequence $\{(s_{t},n_{t})\}_{t=0}^{l+1}$ with $1\leq n_{l+1}<n_{l}<\cdots<n_{1}<n_{0}=j$ until
$\left|{\left(y_{s_{l+1}}^{\left(n_{l+1}\right)}\right)^{\top}R_{j}y_{s_{l+1}}^{\left(n_{l+1}\right)}}\right|<\frac{3}{8}-\frac{\varepsilon_{n_{l+1}}}{2}.$
Inequality (39) holds when the superscript $j=1$ since $T_{1}=[\alpha_{1}],\ y_{1}^{(1)}=1$ and $z^{(1)}_{1}=q_{1}$. Therefore, we can obtain $\{(s_{t},n_{t})\}_{t=0}^{l+1}$ in finite steps, resulting in the following estimate.
@@
Any Rayleigh quotient of $A_{j}^{*}$ is bounded by its eigenvalues [Parlett, 1998], i.e., for any $\nu^{(j)}$,
@@
where $\lambda_{min}^{(j)}$ and $\lambda_{max}^{(j)}$ are the minimum and maximal eigenvalue of $A_{j}^{*}$, respectively. Based on 3.2 and 3.3, we complete the proof. ∎
@@
We detail the proof for Lemma 3.10 in this section, starting with a proof sketch.
@@ G.1 Proof sketch
The proof of Lemma 3.10 is structured by four steps.
Step1: extending $\varepsilon_{j}$ to $\tilde{\varepsilon}_{j}$ in the lemmas given in Appendix F.
@@
In parallel, we can also view the $A_{1},b_{1}$ as reference values. In this way, denote the similar quantities
@@ G.2 Extended lemmas from Appendix F
(Extended version of Lemma F.2) Suppose Assumptions 3.1 to 3.3 hold. The dynamic Lanczos process in Algorithm 2 with normalized $q_{1}$ satisfies
@@ G.3 Proof of Step2
where $T_{j}$ is nonsingular because of Lemma G.2. By (40), (43), and (46), we have
@@
In this way, $\left|\beta_{j+1}e_{j}^{\top}\left(T_{j}\right)^{-1}e_{1}\right|$ can be seen as the residual in the $j$-th step of the classic Lanczos process with the positive-definite matrix $T$ and the iniitial vector $e_{1}$. Since the eigenvalues of $T$ satisfy (49), it follows from the standard convergence property of the Lanczos process [Greenbaum, 1997] that
@@
G.4 Proof of $\mu_{g}-\frac{2\sqrt{3}}{3}\left(j+1\right)^{3}L_{gyy}{\varepsilon}_{j}>0$
In this part, we will give the detailed proof of $\mu_{g}-\frac{2\sqrt{3}}{3}\left(j+1\right)^{3}L_{gyy}{\varepsilon}_{j}>0$. At the same time, we demonstrate that, with appropriate step sizes, the auxiliary variable $v_{k}$ is bounded, thereby showing that the hyper-gradient estimator (5) remains bounded. This ensures the stable behavior of the dynamic Lanczos process,
@@ G.4 Proof of $\mu_{g}-\frac{2\sqrt{3}}{3}\left(j+1\right)^{3}L_{gyy}{\varepsilon}_{j}>0$
Suppose Assumptions 3.1, 3.2, 3.3, 3.6 and 3.9 hold. If within each epoch, we set the step size $\theta\sim\mathcal{O}({1}/{m})$ a constant for $y$, and the step size for $x$ as zero in the first $m_{0}\sim\mathcal{O}(1)$ steps and the others as an appropriate constant $\lambda\sim\mathcal{O}({1}/{m^{4}})$, then for any epoch,
@@
Then, we give a proof by induction. At the beginning of the algorithm, 3.9 reveals
@@
Then by setting the stepsizes
$\theta\leq\frac{\tilde{\varepsilon}}{\mu_{g}L_{gy}m},$ (51)
$\lambda\leq\frac{\sqrt{3}\tilde{\varepsilon}}{4(m+1)^{4}L_{gyy}}\left({1+\theta\frac{L_{gy}L_{gx}}{\mu_{g}}}\right)^{-1}\left(C_{fx}+\frac{1}{\mu_{g}^{2}}\left(3\sqrt{\tilde{\kappa}}C_{r}+C_{fx}\right)\right)^{-1},$
and noticing that
$\left\|{\nabla_{y}g(x_{i+1},y_{i})}\right\|$ $=\left\|{\nabla_{y}g(x_{i+1},y_{i})-\nabla_{y}g(x_{i+1},y^{*}_{i+1})}\right\|$
$\leq L_{gy}\left\|{y_{i}-y^{*}_{i+1}}\right\|$
$\leq L_{gy}\left({\left\|{y_{i}-y^{*}_{i}}\right\|+\left\|{y_{i}^{*}-y^{*}_{i+1}}\right\|}\right)$
$\leq L_{gy}\left\|{y_{i}-y^{*}_{i}}\right\|+\frac{L_{gy}L_{gx}}{\mu_{g}}\left\|{x_{i}-x_{i+1}}\right\|,$
@@
$\tilde{\varepsilon}_{j+1}\leq$ $\ \lambda\sum_{i=1}^{i=j}{\left\|\widetilde{\nabla}\varphi_{i}\right\|}+\theta\sum_{i=1}^{i=j}{\left\|\nabla_{y}g\left(x_{i+1},y_{i}\right)\right\|}$
$\leq$ $\left({1+\theta\frac{L_{gy}L_{gx}}{\mu_{g}}}\right)\lambda\sum_{i=1}^{i=j}{\left\|\widetilde{\nabla}\varphi_{i}\right\|}+\theta L_{gy}\sum_{i=1}^{i=j}\left\|{y_{i}-y^{*}_{i}}\right\|$
$\leq$ $\left({1+\theta\frac{L_{gy}L_{gx}}{\mu_{g}}}\right)\lambda\sum_{i=1}^{i=j}{\left\|\widetilde{\nabla}\varphi_{i}\right\|}+\theta L_{gy}\sum_{i=1}^{i=j}\varepsilon_{i}$
$\leq$ $\left({1+\theta\frac{L_{gy}L_{gx}}{\mu_{g}}}\right)\lambda\sum_{i=1}^{i=j}{\left\|\widetilde{\nabla}\varphi_{i}\right\|}+\theta L_{gy}\frac{\sqrt{3}\mu_{g}m}{4\left(m+1\right)^{3}L_{gyy}}$
$\leq$ $\ \frac{\sqrt{3}\tilde{\varepsilon}}{2(m+1)^{3}L_{gyy}}.$ (52)
@@
Setting
$\lambda\leq\frac{\sqrt{3}\mu_{g}}{8(m+1)^{4}L_{gyy}}\left(C_{fx}+\frac{1}{\mu_{g}^{2}}\left(3\sqrt{\tilde{\kappa}}C_{r}+C_{fx}\right)\right)^{-1}\left(1+\frac{L_{gx}}{\mu_{g}}+\left(1-\theta\mu_{g}\right)^{\frac{1}{2}}\left(\frac{L_{gx}}{\mu_{g}}\right)\right)^{-1}$
yields from 3.9, (54), (57) that
@@
As for the next epoch, denoting $C_{v}=\frac{1}{\mu_{g}}\left(3\sqrt{\tilde{\kappa}}C_{r}+C_{fx}\right)$, we have
@@
by choosing $m,\tilde{\varepsilon}$ such that
@@
Suppose Assumptions 3.1, 3.2, 3.3, 3.6 and 3.9 hold. If within each epoch, we set the step size $\theta\sim\mathcal{O}({1}/{m})$ a constant for $y$ and the step size for $x$ as zero in the first $m_{0}\sim\mathcal{O}(1)$ steps, and the others as an appropriate constant $\lambda\sim\mathcal{O}({1}/{m^{4}})$, we have the following inequality,
@@
Appendix H Proof of The Main Theorem
In this section, we provide proof of the main theorem presented in Section 3.2. Let $A_{k}=\nabla^{2}_{yy}g(x_{k},y_{k})$ and $b_{k}=\nabla_{y}f(x_{k},y_{k})$, and let the reference values be $A_{k}^{*}=\nabla_{yy}^{2}g\left(x_{k},y^{*}_{k}\right)$, $b^{*}_{k}=\nabla_{y}f\left(x_{k},y^{*}_{k}\right)$ and $v_{k}^{*}=(A_{k}^{*})^{-1}b_{k}^{*}$. To begin with, a short proof sketch is provided for guidance, which is structured in four main steps.
Step1: upper-bounding the residual $\left\|{v_{k}-v_{k}^{*}}\right\|$
Appendix F and G lay a foundation to bound the residual term $\left\|{v_{k}-v_{k}^{*}}\right\|$ (as a corollary of Lemma G.5).
Step2: studying the descent property of $\left\|{y_{k}-y_{k}^{*}}\right\|$
Lemma H.1 reveals the descent property of the estimation error for $y^{*}$ as follows,
Step3: Controlling the hyper-gradient estimation error ~$\|\widetilde{\nabla}\varphi(x_{k})-\nabla\varphi(x_{k})\|$
Defining
$\delta_{k}:=\left(\frac{L_{fx}^{2}\mu_{g}^{2}+L_{gxy}^{2}C_{fg}^{2}}{L_{gx}^{2}\mu_{g}^{2}}\right)\left\|y_{k}-y_{k}^{*}\right\|^{2}+\left\|v_{k}-v_{k}^{*}\right\|^{2},$
and incorporating the results from the last two steps, then in Lemma H.2 we can establish the upper bound for ~$\|\widetilde{\nabla}\varphi(x_{k})-\nabla\varphi(x_{k})\|$ and $\delta_{k}$ recursively.
Step4: Assembling the estimations above and achieving the conclusion
Consider the descent property of $\mathcal{L}_{k}:=\varphi_{k}+\delta_{k}$. Substituting the inequalities developed in step1 to step3, and telescoping the index from 00 to $K$ gives the convergence results.
The following lemma displays the descent property of the iterates $\{y_{k}\}$.
@@ Appendix H Proof of The Main Theorem
Suppose Assumptions 3.2 and 3.3 hold. Setting $0<\theta\leq\frac{2}{\mu_{g}+L_{gy}}$, we have
@@
Algorithm 2 executes a single-step gradient descent on the strongly convex function $g(x_{k+1},\cdot)$ during the outer iteration. Leveraging the established convergence properties of strongly convex functions [Nesterov et al., 2018], we are thus able to derive the following.
@@
Suppose Assumptions 3.1, 3.2, 3.3, 3.6 and 3.9 hold. Within each epoch, we set the step size $\theta\sim\mathcal{O}({1}/{m})$ a constant for $y$ and the step size for $x$ as zero in the first $m_{0}$ steps, and the others as an appropriate constant $\lambda\sim\mathcal{O}({1}/{m^{4}})$, then the iterates
@@
~$\left\|\widetilde{\nabla}\varphi\left(x_{k}\right)-\nabla\varphi\left(x_{k}\right)\right\|^{2}\leq$ $~{}3L_{gx}^{2}\delta_{k},$ (63)
and
$\delta_{k}\leq$ $\ \iota^{2\left(m-m_{0}\right)}\left(\delta_{m\left(h-1\right)}\right)+\iota^{2m}\delta_{m\left(h-1\right)}+\iota^{2m}\delta_{m\left(h-2\right)}$
$+12m^{2}\lambda^{2}\omega_{\varphi}\Big{(}\|\widetilde{\nabla}\varphi(x_{mh})\|^{2}+\|\widetilde{\nabla}\varphi(x_{m(h-1)})\|^{2}+\|\widetilde{\nabla}\varphi(x_{m(h-2)})\|^{2}$
$+\sum_{t=m_{0}}^{j-1}{\|\widetilde{\nabla}\varphi(x_{mh+t})\|^{2}}+\sum_{t=m_{0}}^{m-1}{\|\widetilde{\nabla}\varphi(x_{m(h-1)+t})\|^{2}}+\sum_{t=m_{0}}^{m-1}{\|\widetilde{\nabla}\varphi(x_{m(h-2)+t})\|^{2}}\Big{)},$ (64)
where $10<\iota<$ 1, $m_{0}\sim\Omega(\log m)$ are constants, $m$ is the subspace dimension,
$\delta_{k}=\left(\frac{L_{fx}^{2}\mu_{g}^{2}+L_{gxy}^{2}C_{fg}^{2}}{L_{gx}^{2}\mu_{g}^{2}}\right)\left\|y_{k}-y_{k}^{*}\right\|^{2}+\left\|v_{k}-v_{k}^{*}\right\|^{2}.$
@@
By (H), conclusion from Lemma G.3 and the Young’s inequality, it holds that
@@
By definition of $\varepsilon_{j}^{h}$, the update rule of $x$, and Young’s inequality, it holds that
$\left(\varepsilon_{j}^{h}\right)^{2}\leq 2m\lambda^{2}\left(1+\frac{L_{gx}}{\mu_{g}}\right)^{2}\sum_{i=1}^{j-1}{\left\|\widetilde{\nabla}\varphi\left(x_{mh+i}\right)\right\|^{2}}+2\left\|y_{mh+i\left(j\right)}-y_{mh+i\left(j\right)}^{*}\right\|^{2}$
for some $m_{0}+1\leq i(j)\leq j$. Then we apply the descent property of $\|y_{k}-y_{k}^{*}\|$ (Lemma H.1) recursively to derive
$\left(\varepsilon_{j}^{h}\right)^{2}\leq$ $\ 2m\lambda^{2}\left(1+\frac{L_{gx}}{\mu_{g}}\right)^{2}\sum_{i=1}^{j-1}{\left\|\widetilde{\nabla}\varphi\left(x_{mh+i}\right)\right\|^{2}}+2\left(1+\sigma\right)^{i\left(j\right)}\left(1-\theta\mu_{g}\right)^{i\left(j\right)}\left\|y_{mh}-y_{mh}^{*}\right\|^{2}$
$+2\left(1+\frac{1}{\sigma}\right)\left(1-\theta\mu_{g}\right)\left(\frac{L_{gx}}{\mu_{g}}\right)^{2}\sum_{r=0}^{i\left(j\right)}{\left(1+\sigma\right)^{r}\left(1-\theta\mu_{g}\right)^{r}\left\|x_{k-r}-x_{k-r-1}\right\|^{2}}$
$\leq$ $\ 2\lambda^{2}\left(1+\frac{L_{gx}}{\mu_{g}}\right)^{2}\left(m+\left(1+\frac{1}{\sigma}\right)\left(1-\theta\mu_{g}\right)\right)\sum_{i=1}^{j-1}{\left\|\widetilde{\nabla}\varphi\left(x_{mh+i}\right)\right\|^{2}}+2\left\|y_{mh}-y_{mh}^{*}\right\|^{2}$
$\leq$ $\ 2\lambda^{2}\left(1+\frac{L_{gx}}{\mu_{g}}\right)^{2}\left(m+\left(1+\frac{1}{\sigma}\right)\left(1-\theta\mu_{g}\right)\right)\sum_{i=1}^{j-1}{\left\|\widetilde{\nabla}\varphi\left(x_{mh+i}\right)\right\|^{2}}$
$+2\left(1+\sigma\right)^{m_{0}}\left(1-\theta\mu_{g}\right)^{m_{0}}\left\|y_{mh}-y_{mh}^{*}\right\|^{2}.$ (66)
Since Lemma G.4 reveals that under the appropriate step-size setting $\lambda\sim\mathcal{O}(\frac{1}{m^{4}})$ and $\theta\sim\mathcal{O}(\frac{1}{m})$, $\tilde{\kappa}(j)\leq\tilde{\kappa}$ for $\tilde{\kappa}\sim\Omega(\frac{L_{gy}}{\mu_{gy}})$, we also choose $\sigma>0$ such that